OpenStudy (anonymous):

The length of a rectangle is five times its width. If the perimeter is at most 96 centimeters, what is the greatest possible value for the width? A. 40 cm B. 19.2 cm C. 16 cm D. 8 cm

OpenStudy (anonymous):

@Destinymasha @campbell_st @beccaboo333 @Awesome781

OpenStudy (jonnyvonny):

From the question, we are able to derive an equation; do you know what that is? Also, for future use, we have the equation P=2L+2W (from the mathematics chart, or, if you think about it, that is what the perimeter of a rectangle would be [since you're just adding all the sides]).

OpenStudy (reemii):

in the formulas of the area and perimeter, you can use L = 5W, that's what the first tells you. so, "perimeter is at most 96" -> \(L + W + L + W = 5W + W + 5W + W \le 96\) then..?

OpenStudy (reemii):

the first *sentence*

OpenStudy (reemii):

Do you see why \(12W\) is the perimeter? \(12W\le 96 \iff W \le 8\)..